OpenStudy (jasonderubis):

How are impulse and momentum related? Impulse and momentum are the same thing. Momentum is the change in impulse acting on an object. Impulse and momentum are always equal in magnitude. Impulse is the change in momentum of an object.

OpenStudy (shamim):

Accaccording to newton's second law of motion F=(mv-mu)/t

OpenStudy (shamim):

F*t=mv-mu

OpenStudy (shamim):

F*t=impulse

OpenStudy (shamim):

mv-mu=change of momentum of the object